a clamped string is oscillating in nth harmonic, then 
(a) total energy of oscillating will be n^2 times that of fundamental frequency.
(b) total energy of oscillations will be (n-1)^2 times that of fundamental frequency.
© average kinetic energy of the string over a complete oscillations is half of that of the total energy of the string 
(D) none of these 
 
More than one correct type.

Grade:11

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

the total energy of oscillations will be n^2 times that of the fundamental frequency

the average kinetic energy of the string over complete oscillations is half of that of the total energy of the string.

Regards
Arun
4 years ago
For a sine wave, y=Asin(kxΩt)
Velocity equation for this wave is Vy=ΩAcos(kxΩt)
Kinetic energy = d(KE)=1/2(Vy2×dm)=1/2(Vy2×μdx)μ is the linear mass density.

=> 1/2(μ×Ω2×A2×cos2(kxΩt))dx
integrating at t=0, with limits as 0 and λ, we have
 
K.E=1/4(μ×Ω2×A2×λ)
Potential energy, dU=1/2(Ω2×y2×μ)dx

integrating at t=0, with limits as 0 and λ, we have
U=1/4(μ×Ω2×A2×λ)

Total energy E=K.E+U
=> E=1/2(μA2λ)
Therefore, for the first and fundamental frequency, energy is 
E1=(1/2(μA2λ))/n2
And clearly from the above derivation, we have, K.E is half the total energy.